OpenStudy (anonymous):

A coffee shop pays Coffee Delivery Company A, a certain price for each disposable cup it orders plus a weekly delivery charge to remain on the driver’s delivery route. The cups are purchased in increments of 500. To quickly determine how much the coffee shop will be spending on cups before their arrival, the owner created the following table

OpenStudy (anonymous):

Help pleasee! ;((

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Alright so for average rate of change... from 0 to 1000 cups... \[\large \frac{f(1000) - f(0)}{1000 - 0} = \large \frac{44 - 7}{1000} = \frac{37}{1000} = .037\] And then for all 3000 cups \[\large \frac{f(3000) - f(0)}{3000 - 0} = \large \frac{118 - 7}{3000} = \frac{111}{3000} = .037\] So sine they are the same rate of change...we can tell that this is a linear function

OpenStudy (johnweldon1993):

Part 2) If we look at the table...the price at 0 cups...is 7 dollars ...a possible reason for this? well maybe there is just a 7 dollar charge for looking at the cups :D lol

OpenStudy (johnweldon1993):

Part 3) well the equation of this function will be P(x) = .037x + 7 (where 'x' is in every cup purchased) Why? because the average rate of change...is .037 cents per cup :) So the graph would have a y-intercept at 7...and a slope of .037  something like that

OpenStudy (johnweldon1993):

Part D Well we just did that.....the equation would be P(x) = .037x + 7 :)

OpenStudy (johnweldon1993):

Part E ...well we can't answer it....there's no question to answer lol

OpenStudy (anonymous):

i didnt understand B? what would i write?

OpenStudy (johnweldon1993):

Well now that I read E ... p(0) = 7 This can be because the cost per week to have cups delivered is $7 even if there are no cups bought

OpenStudy (anonymous):

Im lost was that for e or b? ^

OpenStudy (johnweldon1993):

'b' :)